Ticket Number: HEA-NOC/20100429-3 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: BT Ireland
Ticket Opened: 20100429 10:22 UTC+1 Problem Start: 20100429 10:00 UTC+1
Ticket Closed: 20100504 09:55 UTC+1 Problem End: 20100429 20:30 UTC+1
Site/line: NCAD, NUIM
Two BT circuits are after going down HEAN105011(NUIM) and HEAN107593(NCAD)
This is causing a full site outage to NCAD and a loss of resilience for
20100429 10:19 TG Opening ticket with BT
20100429 12:41 BN BT have reported a fibre break affecting multiple BT
customers. A crew is on-site, but the repair window is up until 19:00
20100429 16:16 TG Fibre splicing is now scheduled for 18:00 this evening,
we expect this issue to be resolved at 20:00. The issue was caused by a
third party damaging fibre in CWT.
20100429 16:44 EW BT Ireland Change Notification : 100429-INFRA-CAT3
Planned Engineering Works scheduled to commence at 18:00 on 29/04/10. An
emergency planned engineering work is required to repair broken cables in
City West which have been damaged by a third party.
An emergency work is required to repair the damage.
Work will commence at 18:00 and run for approximately 2 hours.
The circuits listed below will be at risk during this work.
HEAN0305 at risk of outage - from UCD-ROADM to CWT-ROADM
HEAN103306 at risk of outage - from Esat-X to NUI Maynooth Gig-e -
"Protected" by HEAN105011
HEAN105011 at risk of outage - from Esat-X to NUI Maynooth - Gig-e -
"Protection" for HEAN103306
HEAN107593 at risk of outage - from Esat-X to NCAD
Also at risk during the work either to a loss of diversity or an outage
HEAN107467 - Loss of diversity – from Esat-X to CDVEC Whitehall
HEAN107482/VC3S.01 - Loss of diversity - from Esat-X to CDVEC Killester
HEAN107482/VC3S.02 - Loss of diversity – from Esat-X to CDVEC Killester
HEAN107482/VCG3 - Loss of diversity – from Esat-X to CDVEC Killester
HEAN107482 - Loss of diversity – from Esat-X to CDVEC Killester
HEAN107499/VC3S.01 - Loss of diversity - from Esat-X to CDVEC Liberties FE
HEAN107499/VC3S.02 - Loss of diversity – from Esat-X to CDVEC Liberties FE
HEAN107499/VCG3.01 - Loss of diversity – from Esat-X to CDVEC Liberties FE
HEAN107499 - Loss of diversity – from Esat-X to CDVEC Liberties FE
HEAN107523 - Loss of diversity – from Tipperary Institute in Thurles to
HEAN109489 - Loss of diversity- from Esat-X to EPA, Johnstown
20100430 09:36 PMD BT restored connectivity at 20:30 last night. The links
have been up and stable since then. We will keep monitoring the link for a
the time being but no further issues are expected.
20100430 09:54 PMD taking ticket ownership
20100430 18:17 PMD The connection has been stable for almost 22 hours.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20100429-3
HEAnet Limited. Registered in Ireland, No. 275301.